# Low-metallicity massive single stars with rotation

### v2

@@ Acknowledgements.
We thank both our anonymous referees, the first for making us add the entirety of Sect. 4 on comparing our results with all available observations from the literature, which made this work a truly unique contribution to the field, and the second for spotting mistakes and raising our attention to various omissions, such as the importance of including NE models when computing the spectral hardness. This research was funded in part by the National Science Center (NCN), Poland under grant number OPUS 2021/41/B/ST9/00757. For the purpose of Open Access, the author has applied a CC-BY public copyright license to any Author Accepted Manuscript (AAM) version arising from this submission. D.Sz. was supported by the Alexander von Humboldt Foundation. AACS is funded by the German Deutsche Forschungsgemeinschaft, DFG in the form of an Emmy Noether Research Group – Project-ID 445674056 (SA4064/1-1, PI Sander). AACS further acknowledges financial support by the Federal Ministry for Economic Affairs and Climate Action (BMWK) via the German Aerospace Center (Deutsches Zentrum für Luft- und Raumfahrt, DLR) grant 50 OR 2503 (PI: Sander). This project was co-funded by the European Union (Project 101183150 - OCEANS). This research is partially done in the frame of the European Union’s Framework Programme for Research and Innovation Horizon 2020 (2014-2020) under the Marie Skłodowska-Curie Grant Agreement No. 823734. B.K. and J.K. acknowledge the support of the Grant Agency of the Czech Republic (GAČR 25-15910S) and RVO:67985815. CK acknowledges financial support from the State Agency for Research of the Spanish MCIU through Center of Excellence Severo Ochoa award to the Instituto de Astrofísica de Andalucía CEX2021- 001131-S funded by MCIN/AEI/10.13039/501100011033, and from the grant PID2022-136598NB-C32 “Estallidos8”. M. Garcia gratefully acknowledges support by grants PID2019-105552RB-C41 and PID2022-137779OB-C41, funded by the Spanish Ministry of Science, Innovation and Universities/State Agency of Research MICIU/AEI/10.13039/501100011033 and by “ERDF A way of making Europe”. D. Sz. is grateful for the relevant discussions with S. Heap, R. Sarwar, H. Stinshoff, K. Sen, G. Telford, N. Langer, P. Crowther, C. Leitherer, G. Gräfener, I. Mandel, S. Justham, C. Neijssel, A. Vigna-Gómez, and S. Vinciguerra, and Á. Szabó.References
- Aadland et al. (2022) Aadland, E., Massey, P., Hillier, D. J., et al. 2022, ApJ, 931, 157
- Abdellaoui (2023) Abdellaoui, S. 2023, PhD thesis, MASARYK UNIVERSITY, Faculty of Science
- Agrawal et al. (2022) Agrawal, P., Szécsi, D., Stevenson, S., Eldridge, J. J., & Hurley, J. 2022, MNRAS, 512, 5717
- Aguilera-Dena et al. (2018) Aguilera-Dena, D. R., Langer, N., Moriya, T. J., & Schootemeijer, A. 2018, ApJ, 858, 115
- Aloisi et al. (1999) Aloisi, A., Tosi, M., & Greggio, L. 1999, ApJ, 118, 302
- Annibali et al. (2013) Annibali, F., Cignoni, M., Tosi, M., et al. 2013, AJ, 146, 144
- Arroyo-Polonio et al. (2023) Arroyo-Polonio, A., Iglesias-Páramo, J., Kehrig, C., et al. 2023, A&A, 677, A114
- Arroyo-Polonio et al. (2024) Arroyo-Polonio, A., Kehrig, C., Iglesias-Páramo, J., et al. 2024, A&A, 687, A77
- Barlow & Hummer (1982) Barlow, M. J. & Hummer, D. G. 1982, in IAU Symposium, Vol. 99, Wolf-Rayet Stars: Observations, Physics, Evolution, ed. C. W. H. De Loore & A. J. Willis, 387–392
- Beals (1933) Beals, C. S. 1933, The Observatory, 56, 196
- Berg et al. (2022) Berg, D. A., James, B. L., King, T., et al. 2022, ApJS, 261, 31
- Bestenlehner et al. (2020) Bestenlehner, J. M., Crowther, P. A., Caballero-Nieves, S. M., et al. 2020, MNRAS, 499, 1918
- Bowler et al. (2017) Bowler, R. A. A., McLure, R. J., Dunlop, J. S., et al. 2017, MNRAS, 469, 448
- Brott et al. (2011) Brott, I., de Mink, S. E., Cantiello, M., et al. 2011, A&A, 530, A115
- Brown et al. (2002) Brown, T. M., Heap, S. R., Hubeny, I., Lanz, T., & Lindler, D. 2002, ApJ, 579, L75
- Burssens et al. (2023) Burssens, S., Bowman, D. M., Michielsen, M., et al. 2023, Nature Astronomy, 7, 913
- Cannon & Pickering (1916) Cannon, A. J. & Pickering, E. C. 1916, Annals of Harvard College Observatory, 76, 19
- Christensen et al. (2016) Christensen, C. R., Davé, R., Governato, F., et al. 2016, ApJ, 824, 57
- Chruslinska & Nelemans (2019) Chruslinska, M. & Nelemans, G. 2019, MNRAS, 488, 5300
- Crowther et al. (1998) Crowther, P. A., De Marco, O., & Barlow, M. J. 1998, MNRAS, 296, 367
- Crowther & Hadfield (2006) Crowther, P. A. & Hadfield, L. J. 2006, A&A, 449, 711
- Crowther et al. (1995) Crowther, P. A., Hillier, D. J., & Smith, L. J. 1995, A&A, 293, 172
- Crowther & Walborn (2011) Crowther, P. A. & Walborn, N. R. 2011, MNRAS, 416, 1311
- de Mink et al. (2009) de Mink, S. E., Cantiello, M., Langer, N., et al. 2009, A&A, 497, 243
- de Mink et al. (2013) de Mink, S. E., Langer, N., Izzard, R. G., Sana, H., & de Koter, A. 2013, ApJ, 764, 166
- de Mink & Mandel (2016) de Mink, S. E. & Mandel, I. 2016, MNRAS, 460, 3545
- Dorozsmai et al. (2024) Dorozsmai, A., Toonen, S., Vigna-Gómez, A., de Mink, S. E., & Kummer, F. 2024, MNRAS, 527, 9782
- du Buisson et al. (2020) du Buisson, L., Marchant, P., Podsiadlowski, P., et al. 2020, MNRAS, 499, 5941
- Edlen (1932) Edlen, B. 1932, The Observatory, 55, 115
- Eldridge & Stanway (2012) Eldridge, J. J. & Stanway, E. R. 2012, MNRAS, 419, 479
- Eldridge & Stanway (2016) Eldridge, J. J. & Stanway, E. R. 2016, MNRAS, 462, 3302
- Eldridge et al. (2019) Eldridge, J. J., Stanway, E. R., & Tang, P. N. 2019, MNRAS, 482, 870
- Evans et al. (2019) Evans, C. J., Castro, N., Gonzalez, O. A., et al. 2019, A&A, 622, A129
- Franeck et al. (2022) Franeck, A., Wünsch, R., Martínez-González, S., et al. 2022, ApJ, 927, 212
- Garcia et al. (2019) Garcia, M., Herrero, A., Najarro, F., Camacho, I., & Lorenzo, M. 2019, MNRAS, 484, 422
- Georgy et al. (2012) Georgy, C., Ekström, S., Meynet, G., et al. 2012, A&A, 542, A29
- Girichidis et al. (2018) Girichidis, P., Seifried, D., Naab, T., et al. 2018, MNRAS, 480, 3511
- Götberg et al. (2017) Götberg, Y., de Mink, S. E., & Groh, J. H. 2017, A&A, 608, A11
- Gräfener et al. (2002) Gräfener, G., Koesterke, L., & Hamann, W.-R. 2002, A&A, 387, 244
- Groh et al. (2019) Groh, J. H., Ekström, S., Georgy, C., et al. 2019, A&A, 627, A24
- Groh et al. (2014) Groh, J. H., Meynet, G., Ekström, S., & Georgy, C. 2014, A&A, 564, A30
- Gull et al. (2022) Gull, M., Weisz, D. R., Senchyna, P., et al. 2022, ApJ, 941, 206
- Guseva et al. (2000) Guseva, N. G., Izotov, Y. I., & Thuan, T. X. 2000, ApJ, 531, 776
- Haid et al. (2019) Haid, S., Walch, S., Seifried, D., et al. 2019, MNRAS, 482, 4062
- Hainich et al. (2018) Hainich, R., Oskinova, L. M., Shenar, T., et al. 2018, A&A, 609, A94
- Hainich et al. (2015) Hainich, R., Pasemann, D., Todt, H., et al. 2015, A&A, 581, A21
- Hamann & Gräfener (2003) Hamann, W.-R. & Gräfener, G. 2003, A&A, 410, 993
- Hamann & Gräfener (2004) Hamann, W.-R. & Gräfener, G. 2004, A&A, 427, 697
- Hamann et al. (1995) Hamann, W.-R., Koesterke, L., & Wessolowski, U. 1995, A&A, 299, 151
- Hamann et al. (2015) Hamann, W. R., Sander, A., & Todt, H. 2015, in Wolf-Rayet Stars: Proceedings of an International Workshop held in Potsdam, Germany, ed. W.-R. Hamann, A. Sander, & H. Todt
- Heap et al. (2015) Heap, S., Bouret, J.-C., & Hubeny, I. 2015, ArXiv e-print 1504.02742
- Higgins & Vink (2019) Higgins, E. R. & Vink, J. S. 2019, A&A, 622, A50
- Hirschauer et al. (2024) Hirschauer, A. S., Crouzet, N., Habel, N., et al. 2024, AJ, 168, 23
- Hopkins et al. (2011) Hopkins, P. F., Quataert, E., & Murray, N. 2011, MNRAS, 417, 950
- Hunter & Thronson (1995) Hunter, D. & Thronson, H. 1995, ApJ, 452, 238
- Izotov & Thuan (2002) Izotov, Y. & Thuan, T. 2002, ApJ, 567
- Izotov & Thuan (2004) Izotov, Y. & Thuan, T. 2004, ApJ, 616, 768
- Izotov et al. (1997) Izotov, Y. I., Foltz, C. B., Green, R. F., Guseva, N. G., & Thuan, T. X. 1997, ApJ, 487, L37
- Izotov et al. (2011) Izotov, Y. I., Guseva, N. G., & Thuan, T. X. 2011, ApJ, 728, 161
- Izotov et al. (2016) Izotov, Y. I., Schaerer, D., Thuan, T. X., et al. 2016, MNRAS, 461, 3683
- Izotov & Thuan (1998) Izotov, Y. I. & Thuan, T. X. 1998, ApJ, 497, 227
- Jaskot & Oey (2013) Jaskot, A. E. & Oey, M. S. 2013, ApJ, 766, 91
- Kehrig et al. (2021) Kehrig, C., Guerrero, M. A., Vílchez, J. M., & Ramos-Larios, G. 2021, ApJ, 908, L54
- Kehrig et al. (2013) Kehrig, C., Pérez-Montero, E., Vílchez, J. M., et al. 2013, MNRAS, 432, 2731
- Kehrig et al. (2018) Kehrig, C., Vílchez, J. M., Guerrero, M. A., et al. 2018, MNRAS, 480, 1081
- Kehrig et al. (2015a) Kehrig, C., Vílchez, J. M., Pérez-Montero, E., et al. 2015a, in Wolf-Rayet Stars: Proceedings of an International Workshop held in Potsdam, Germany, 1-5 June 2015. Edited by Wolf-Rainer Hamann, Andreas Sander, Helge Todt. Universitätsverlag Potsdam, 2015., p.55-58, ed. W.-R. Hamann, A. Sander, & H. Todt, 55–58
- Kehrig et al. (2015b) Kehrig, C., Vílchez, J. M., Pérez-Montero, E., et al. 2015b, ApJ, 801, L28
- Kehrig et al. (2016) Kehrig, C., Vílchez, J. M., Pérez-Montero, E., et al. 2016, MNRAS, 459, 2992
- Klessen & Glover (2023) Klessen, R. S. & Glover, S. C. O. 2023, ARA&A, 61, 65
- Kubátová et al. (2019) Kubátová, B., Szécsi, D., Sander, A. A. C., et al. 2019, A&A, 623, A8
- Kunth & Östlin (2000) Kunth, D. & Östlin, G. 2000, A&A Rev., 10, 1
- Lahén et al. (2023) Lahén, N., Naab, T., Kauffmann, G., et al. 2023, MNRAS, 522, 3092
- Langer (1989) Langer, N. 1989, A&A, 210, 93
- Lebouteiller et al. (2013) Lebouteiller, V., Heap, S., Hubeny, I., & Kunth, D. 2013, A&A, 553, A16
- Lebouteiller et al. (2017) Lebouteiller, V., Péquignot, D., Cormier, D., et al. 2017, A&A, 602, A45
- Lecavelier des Etangs et al. (2004) Lecavelier des Etangs, A., Désert, J.-M., Kunth, D., et al. 2004, A&A, 413, 131
- Lecroq et al. (2024) Lecroq, M., Charlot, S., Bressan, A., et al. 2024, MNRAS, 527, 9480
- Liu et al. (2025) Liu, B., Sibony, Y., Meynet, G., & Bromm, V. 2025, ApJ, 980, L30
- López-Sánchez & Esteban (2010a) López-Sánchez, Á. R. & Esteban, C. 2010a, A&A, 516, A104
- López-Sánchez & Esteban (2010b) López-Sánchez, Á. R. & Esteban, C. 2010b, A&A, 517, A85
- Maeder & Meynet (2000) Maeder, A. & Meynet, G. 2000, ARA&A, 38, 143
- Mandel & de Mink (2016) Mandel, I. & de Mink, S. E. 2016, MNRAS, 458, 2634
- Marchant et al. (2017) Marchant, P., Langer, N., Podsiadlowski, P., et al. 2017, A&A, 604, A55
- Marchant et al. (2016) Marchant, P., Langer, N., Podsiadlowski, P., Tauris, T. M., & Moriya, T. J. 2016, Astronomy & Astrophysics, 588, A50
- Meynet & Maeder (2000) Meynet, G. & Maeder, A. 2000, A&A, 361, 101
- Micheva et al. (2017) Micheva, G., Oey, M. S., Jaskot, A. E., & James, B. L. 2017, ApJ, 845, 165
- Mintz et al. (2025) Mintz, A., Telford, O. G., Kirby, E. N., et al. 2025, arXiv e-prints, arXiv:2503.19020
- Mokiem et al. (2006) Mokiem, M. R., de Koter, A., Evans, C. J., et al. 2006, A&A, 456, 1131
- Nugis & Lamers (2000) Nugis, T. & Lamers, H. 2000, A&A, 360, 227
- Orlitová et al. (2018) Orlitová, I., Verhamme, A., Henry, A., et al. 2018, A&A, 616, A60
- Oskinova & Schaerer (2022) Oskinova, L. M. & Schaerer, D. 2022, A&A, 661, A67
- Papaderos et al. (2002) Papaderos, P., Izotov, Y. I., Thuan, T. X., et al. 2002, A&A, 393, 461
- Papaderos & Östlin (2012) Papaderos, P. & Östlin, G. 2012, A&A, 537, A126
- Payne (1930) Payne, C. H. 1930, Harvard College Observatory Bulletin, 878, 1
- Péquignot (2008) Péquignot, D. 2008, A&A, 478, 371
- Perley et al. (2016) Perley, D. A., Tanvir, N. R., Hjorth, J., et al. 2016, ApJ, 817, 8
- Riley et al. (2021) Riley, J., Mandel, I., Marchant, P., et al. 2021, MNRAS, 505, 663
- Romagnolo et al. (2022) Romagnolo, A., Belczynski, K., Klencki, J., et al. 2022, arXiv e-prints, arXiv:2211.15800
- Roy et al. (2025) Roy, A., Krumholz, M. R., Salvadori, S., et al. 2025, A&A, 696, A29
- Salpeter (1955) Salpeter, E. 1955, ApJ, 121, 161
- Sana et al. (2012) Sana, H., de Mink, S. E., de Koter, A., et al. 2012, Science, 337, 444
- Sander et al. (2015) Sander, A., Shenar, T., Hainich, R., et al. 2015, A&A, 577, A13
- Sander (2022) Sander, A. A. C. 2022, arXiv e-prints, arXiv:2211.05424
- Sander & Vink (2020) Sander, A. A. C. & Vink, J. S. 2020, MNRAS, 499, 873
- Schaerer et al. (2019) Schaerer, D., Fragos, T., & Izotov, Y. I. 2019, A&A, 622, L10
- Schootemeijer & Langer (2018) Schootemeijer, A. & Langer, N. 2018, A&A, 611, A75
- Searle & Sargent (1972) Searle, L. & Sargent, W. 1972, ApJ, 173, 25
- Seifried et al. (2017) Seifried, D., Walch, S., Girichidis, P., et al. 2017, MNRAS, 472, 4797
- Sen et al. (2024) Sen, K., El Mellah, I., Langer, N., et al. 2024, arXiv e-prints, arXiv:2406.08596
- Senchyna et al. (2019) Senchyna, P., Stark, D. P., Chevallard, J., et al. 2019, arXiv e-prints [arXiv:1904.01615]
- Senchyna et al. (2020) Senchyna, P., Stark, D. P., Mirocha, J., et al. 2020, MNRAS, 494, 941
- Sharpe et al. (2024) Sharpe, K., van Son, L. A. C., de Mink, S. E., et al. 2024, ApJ, 966, 9
- Shirazi & Brinchmann (2012) Shirazi, M. & Brinchmann, J. 2012, MNRAS, 421, 1043
- Smith et al. (1996) Smith, L. F., Shara, M. M., & Moffat, A. F. J. 1996, MNRAS, 281, 163
- Smith et al. (2023) Smith, L. J., Oey, M. S., Hernandez, S., et al. 2023, ApJ, 958, 194
- Sobral et al. (2019) Sobral, D., Matthee, J., Brammer, G., et al. 2019, MNRAS, 482, 2422
- Sobral et al. (2015) Sobral, D., Matthee, J., Darvish, B., et al. 2015, ApJ, 808, 139
- Spencer et al. (2018) Spencer, M. E., Mateo, M., Olszewski, E. W., et al. 2018, AJ, 156, 257
- Stanway & Eldridge (2019) Stanway, E. R. & Eldridge, J. J. 2019, A&A, 621, A105
- Stanway et al. (2016) Stanway, E. R., Eldridge, J. J., & Becker, G. D. 2016, MNRAS, 456, 485
- Stasińska et al. (2015) Stasińska, G., Izotov, Y., Morisset, C., & Guseva, N. 2015, A&A, 576, A83
- Stevenson et al. (2019) Stevenson, S., Sampson, M., Powell, J., et al. 2019, ApJ, 882, 121
- Szécsi (2016) Szécsi, D. 2016, PhD thesis, Mathematisch-Naturwissenschaftlichen Fakultät der Universität Bonn
- Szécsi (2017a) Szécsi, D. 2017a, Contributions of the Astronomical Observatory Skalnate Pleso, 47, 108
- Szécsi (2017b) Szécsi, D. 2017b, Proceedings of Science, PoS(MULTIF2017)065, 2017 [arXiv:1710.05655] [arXiv:1710.05655]
- Szécsi et al. (2022) Szécsi, D., Agrawal, P., Wünsch, R., & Langer, N. 2022, A&A, 658, A125
- Szécsi et al. (2015) Szécsi, D., Langer, N., Yoon, S.-C., et al. 2015, A&A, 581, A15
- Szécsi & Wünsch (2019) Szécsi, D. & Wünsch, R. 2019, ApJ, Vol. 871, nr. 1 [arXiv:1809.01395]
- Thuan & Izotov (2005) Thuan, T. X. & Izotov, Y. I. 2005, ApJS, 161, 240
- Tramper et al. (2013) Tramper, F., Gräfener, G., Hartoog, O. E., et al. 2013, A&A, 559, A72
- Šurlan et al. (2013) Šurlan, B., Hamann, W.-R., Aret, A., et al. 2013, A&A, 559, A130
- Vaduvescu et al. (2007) Vaduvescu, O., McCall, M. L., & Richer, M. G. 2007, AJ, 134, 604
- Vigna-Gómez et al. (2025) Vigna-Gómez, A., Grishin, E., Stegmann, J., et al. 2025, arXiv e-prints, arXiv:2503.17006
- Vigna-Gómez et al. (2018) Vigna-Gómez, A., Neijssel, C. J., Stevenson, S., et al. 2018, MNRAS, 481, 4009
- Vílchez & Iglesias-Páramo (1998) Vílchez, J. M. & Iglesias-Páramo, J. 1998, ApJ, 508, 248
- Vink et al. (2001) Vink, J., de Koter, A., & Lamers, H. 2001, A&A, 369, 574
- Visbal et al. (2016) Visbal, E., Haiman, Z., & Bryan, G. L. 2016, MNRAS, 460, L59
- Weisz et al. (2014) Weisz, D. R., Dolphin, A. E., Skillman, E. D., et al. 2014, ApJ, 789, 147
- Woosley & Heger (2006) Woosley, S. E. & Heger, A. 2006, ApJ, 637, 914
- Yang et al. (2016) Yang, H., Malhotra, S., Gronke, M., et al. 2016, ApJ, 820, 130
- Yoon et al. (2012) Yoon, S.-C., Dierks, A., & Langer, N. 2012, A&A, 542, A113
- Yoon & Langer (2005) Yoon, S.-C. & Langer, N. 2005, A&A, 443, 643
- Yoon et al. (2006) Yoon, S.-C., Langer, N., & Norman, C. 2006, A&A, 460, 199
- Zhao et al. (2013) Zhao, Y., Gao, Y., & Gu, Q. 2013, ApJ, 764, 44
Appendix A Line luminosities
@@ Appendix B Ionizing flux predicted by our PoWR stellar atmosphere models
Models with reduced mass loss, cf. Paper II (for reference only, not used in population synthesis) | |||||||||||

### v3

@@ Acknowledgements.
We thank both our anonymous referees, the first for making us add the entirety of Sect. 4 on comparing our results with all available observations from the literature, which made this work a truly unique contribution to the field, and the second for spotting mistakes and raising our attention to various omissions, such as the importance of including NE models when computing the spectral hardness. This research was funded in part by the National Science Center (NCN), Poland under grant number OPUS 2021/41/B/ST9/00757. For the purpose of Open Access, the author has applied a CC-BY public copyright license to any Author Accepted Manuscript (AAM) version arising from this submission. D.Sz. was supported by the Alexander von Humboldt Foundation. AACS is funded by the German Deutsche Forschungsgemeinschaft, DFG in the form of an Emmy Noether Research Group – Project-ID 445674056 (SA4064/1-1, PI Sander). AACS further acknowledges financial support by the Federal Ministry for Economic Affairs and Climate Action (BMWK) via the German Aerospace Center (Deutsches Zentrum für Luft- und Raumfahrt, DLR) grant 50 OR 2503 (PI: Sander). This project was co-funded by the European Union (Project 101183150 - OCEANS). This research is partially done in the frame of the European Union’s Framework Programme for Research and Innovation Horizon 2020 (2014-2020) under the Marie Skłodowska-Curie Grant Agreement No. 823734. B.K. and J.K. acknowledge the support of the Grant Agency of the Czech Republic (GAČR 25-15910S) and RVO:67985815. CK acknowledges financial support from the State Agency for Research of the Spanish MCIU through Center of Excellence Severo Ochoa award to the Instituto de Astrofísica de Andalucía CEX2021- 001131-S funded by MCIN/AEI/10.13039/501100011033, and from the grant PID2022-136598NB-C32 “Estallidos8”. M. Garcia gratefully acknowledges support by grants PID2019-105552RB-C41 and PID2022-137779OB-C41, funded by the Spanish Ministry of Science, Innovation and Universities/State Agency of Research MICIU/AEI/10.13039/501100011033 and by “ERDF A way of making Europe”. D. Sz. is grateful for the relevant discussions with S. Heap, R. Sarwar, H. Stinshoff, K. Sen, G. Telford, N. Langer, P. Crowther, C. Leitherer, G. Gräfener, I. Mandel, S. Justham, C. Neijssel, A. Vigna-Gómez, and S. Vinciguerra, and Á. Szabó.Appendix A Line luminosities
@@ Appendix B Ionizing flux predicted by our PoWR stellar atmosphere models
| Models with reduced mass loss, see Kubatova:2019 (for reference only, not used in population synthesis) | |||||||||||